OpenStudy (anonymous):

9. For problems 9 - 11, find the dot product if u = <-1, 3>, v = <2, 4>, and w = <2, -5>. u • v A. 2 B. 8 C.-10 D. 10 E. 4 F. 3 10. u • w A. -11 B. 0 C.1 D. -17 E. 11 F. -1 11. (u + v) • w A. 37 B. -37 C.-33 D. -29 E. 29 F. 9

OpenStudy (raden):

u.v = -1*2 + 3*4 = .... u.w = -1*2 + 3*(-5) = ... (u+v).w = (-1+2)*2 + (3+4)*(-5) = ....